The sum of four consecutive even integers is 132. Find the integers
Answer by ikleyn(52921)

.

Let n be the smallest of the four integers.


Then the integers are  n, n+2. n+4, n+6.


The equation for the total sum is


    n + (n+2) + (n+4) + (n+6) = 132.


Simplify and find n


    4n + 12 = 132

    4n      = 132 - 12

    4n      = 120

     n      = 120/4 = 30.


ANSWER.  The numbers are  30, 32, 34 and 36.

Solved.